### v1

@@ V Advantages and Limitations of TL in KC
TL can have several advantages in the context of KC: Securing annotated data in the healthcare sector is a significant challenge, often requiring considerable effort. TL plays a pivotal role in this context by facilitating the use of existing pre-trained models. These models, which have been educated on large and universally accessible datasets, can be adapted to develop specialized models for diagnosing specific conditions like KC. The advantages of applying pre-trained models are manifold. Firstly, these models, having been exposed to vast amounts of data, have the capability to significantly enhance performance on smaller, task-specific datasets, particularly in diagnosing KC. This approach not only leads to considerable savings in time and financial resources but also accelerates the training process. This is because fine-tuning a pre-trained model for a new task is much faster than starting training from scratch. Furthermore, TL reduces the need for extensive data and computational power for training models on specific tasks. In the realm of medical imaging, where missing or scarce data can pose significant challenges, especially for rare diseases like KC, TL offers a valuable solution by employing models that have been pre-trained on extensive datasets. This method effectively bridges the data gap, providing a way to address the issues of scarce data and domain shift that are prevalent due to the high variability in medical imaging data, making it especially beneficial for scenarios involving small datasets.
@@
The application of TL in diagnosing KC presents several challenges and considerations. One primary challenge is domain adaptation, which necessitates identifying a pre-trained model trained on a domain or task closely related to KC diagnosis. The effectiveness of TL hinges on the alignment between the original and target domains and tasks, with discrepancies potentially undermining the model’s ability to transfer knowledge effectively. Moreover, the risk of overfitting is heightened when fine-tuning pre-trained models on limited datasets, leading to models that perform poorly on new or unseen data. Fine-tuning entails careful adjustments of model layers, learning rates, and regularization methods to avoid such overfitting. Data scarcity for the new task can further complicate fine-tuning, potentially resulting in suboptimal performance. Additionally, TL can be computationally demanding, especially with large pre-trained models, posing scalability challenges. Interpretability issues also arise, as understanding the decision-making process of a model trained on a different task can be complex. The biases in the datasets used for training pre-trained models may carry over, affecting performance on certain data subsets. Moreover, these models, being trained on specific datasets, might not generalize well across different datasets, especially when data distributions vary significantly. Privacy concerns are another critical consideration, particularly when sensitive medical data must be shared with third parties for model training. The architecture of the pre-trained model also plays a crucial role in the success of TL ; an ill-suited architecture can hinder effective fine-tuning. Lastly, the substantial computational resources required for TL can make the use of pre-trained models costly, especially when employing multiple models or conducting several rounds of fine-tuning. These challenges underscore the complexity and nuanced considerations involved in applying TL for medical diagnosis tasks such as KC. Table. VI, provides an overview of the benefits and drawbacks of TL-based kidney tumor techniques.
TABLE VI: Summary of advantages and limitations of TL-based kidney tumor techniques.
Work Year LM MPS Dim Metrics Advantages/Limitations
[136] 2022 CNN+TL Segmentation and classification 2D DC, ACC, Pr, SEN, F1, MSE Noise problem, the substantial size of WSIs and the computational costs
[137] 2022 CNN+TL Segmentation and classification 2D Pr, ACC, F1 The anticipated time was higher
[138] 2022 CNN+TL Segmentation and classification 2D ACC, SEN, SPE, SEN, F1, MSE Not accurate at higher levels and insufficient data
[139] 2022 SHUW+CNN Segmentation and classification 2D SEN, DC, SD Application to only one area and acceptable performance
[140] 2022 ResUNet and DenseUNet Segmentation and classification 2D DC Limited testing size, instability of data distribution
[141] 2022 CNN+TL Segmentation, and classification 3D DC, JI, FN, VS, ROC-AUC High complexity, the field of view not be sufficient
[142] 2022 Histo-Cloud+CNN Segmentation 2D Kappa, TPR, MCC The computational complexity
[143] 2022 TL+AE Segmentation, sample and feature space augmentation 2D ROC-AUC Latent representations, small data volume, imbalance datasets, improved predictive model
[144] 2021 U-Net Segmentation and classification 2D CrM, Pr, SEN, ACC, JI, DC Complex deep network and unnecessary computations have been reduced, low processing speed
[145] 2021 SVM Semantic segmentation, feature extraction and classification 2D ACC, SEN, Pr, F1 Motion artifacts, accuracy issue, limited datasets and
[146] 2020 TL+FCNN Segmentation 2D DC, F1 Reducing in the training time, segmentation accuracy and smaller datasets
Abbreviation: Learning model (LM), Main processing steps (MPS), Dimention (Dim),
VI Key Challenges
@@
VI-A Availability, Adapting and Adjusting Models
@@
VI-B Data Scarcity
Data scarcity presents a significant challenge in the development and efficacy of kidney cancer diagnosis systems, impacting everything from image analysis to the training of deep learning models. Several studies have explored these challenges and proposed methods to mitigate the impact of limited data availability [150]. Peng et al. [151] explore how downsampling and compression affect renal image analysis. They found that data scarcity impacts diagnostic accuracy, highlighting the need for robust data handling techniques to ensure the integrity and usefulness of renal images in diagnosis. Similarly, Guo et al. [152] address the limited training data for kidney segmentation by using a cascaded convolutional neural network, demonstrating that innovative network architectures can partly overcome the challenges posed by sparse data.
In the context of clinical applications, Heller et al. [153] provide the kits19 dataset, which includes 300 kidney tumor cases with CT semantic segmentations and surgical outcomes, aiming to enhance research in kidney cancer by making rich datasets available. This effort underscores the critical role of accessible high-quality datasets in advancing early tumor detection and improving surgical outcomes. Flitcroft et al. [154] and Nassour et al. [155] both discuss the scarcity of reliable diagnostic biomarkers for kidney cancer, particularly in non-invasive strategies like urinary protein analysis and the evaluation of Lynch syndrome’s impact on kidney cancer. These studies highlight the gap in epidemiological data and the need for comprehensive datasets to develop effective diagnostic tools.
Technological advancements in handling data scarcity are also evident in deep learning applications. Alzu’bi et al. [156] and Bhattacharjee et al. [157] discuss the use of CNNs for kidney tumor classification and detection. They focus on optimizing deep learning approaches to perform effectively even with limited data, a crucial advancement given the rarity and variability of KC data. Lastly, Wang et al. [158] delve into the utilization of deep learning techniques for the imaging diagnosis of renal cell carcinoma. They emphasize how these advanced computational tools allow clinicians to diagnose and evaluate renal tumors more accurately and swiftly. One of the key strategies discussed is leveraging unlabeled data, which is a valuable solution for coping with data scarcity. By utilizing semi-supervised learning techniques, the potential of existing datasets can be maximized without the need for extensive labeled data.
VI-C Data Bias and Generalization
@@
VI-D Privacy Concerns, Expertise Shortage and Accuracy Instability
@@
VI-E Knowledge Gain and Data Annotation
@@
VI-F Negative Transfer and Overfitting

### v2

@@ 6 Limitations and key challenges
TL can have several advantages in the context of KC, including securing annotated data in the healthcare sector is a significant challenge, often requiring considerable effort. TL plays a pivotal role in this context by facilitating the use of existing pre-trained models. These models, which have been educated on large and universally accessible datasets, can be adapted to develop specialized models for diagnosing specific conditions like KC. The advantages of applying pre-trained models are manifold. Firstly, these models, having been exposed to vast amounts of data, have the capability to significantly enhance performance on smaller, task-specific datasets, particularly in diagnosing KC. This approach not only leads to considerable savings in time and financial resources but also accelerates the training process. This is because fine-tuning a pre-trained model for a new task is much faster than starting training from scratch. Furthermore, TL reduces the need for extensive data and computational power for training models on specific tasks. In the realm of medical imaging, where missing or scarce data can pose significant challenges, especially for rare diseases like KC, TL offers a valuable solution by employing models that have been pre-trained on extensive datasets. This method effectively bridges the data gap, providing a way to address the issues of scarce data and domain shift that are prevalent due to the high variability in medical imaging data, making it especially beneficial for scenarios involving small datasets.
@@
The application of TL in diagnosing KC presents several challenges and considerations. One primary challenge is DA, which necessitates identifying a pre-trained model trained on a domain or task closely related to KC diagnosis. The effectiveness of TL hinges on the alignment between the original and target domains and tasks, with discrepancies potentially undermining the model’s ability to transfer knowledge effectively. Moreover, the risk of overfitting is heightened when fine-tuning pre-trained models on limited datasets, leading to models that perform poorly on new or unseen data. Fine-tuning entails careful adjustments of model layers, learning rates, and regularization methods to avoid such overfitting. Data scarcity for the new task can further complicate fine-tuning, potentially resulting in suboptimal performance. Additionally, TL can be computationally demanding, especially with large pre-trained models, posing scalability challenges. Interpretability issues also arise, as understanding the decision-making process of a model trained on a different task can be complex. The biases in the datasets used for training pre-trained models may carry over, affecting performance on certain data subsets. Moreover, these models, being trained on specific datasets, might not generalize well across different datasets, especially when data distributions vary significantly. Privacy concerns are another critical consideration, particularly when sensitive medical data must be shared with third parties for model training. The architecture of the pre-trained model also plays a crucial role in the success of TL ; an ill-suited architecture can hinder effective fine-tuning. Lastly, the substantial computational resources required for TL can make the use of pre-trained models costly, especially when employing multiple models or conducting several rounds of fine-tuning. These challenges underscore the complexity and nuanced considerations involved in applying TL for medical diagnosis tasks such as KC.
@@
6.1 Availability, adapting and adjusting models
@@
6.2 Data scarcity
Data scarcity presents a significant challenge in the development and efficacy of KC diagnosis systems, impacting everything from image analysis to the training of DL models. Several studies have explored these challenges and proposed methods to mitigate the impact of limited data availability [150]. Peng et al. [151] explore how downsampling and compression affect renal image analysis. They found that data scarcity impacts diagnostic accuracy, highlighting the need for robust data handling techniques to ensure the integrity and usefulness of renal images in diagnosis. Similarly, Guo et al. [152] address the limited training data for kidney segmentation by using a cascaded CNN, demonstrating that innovative network architectures can partly overcome the challenges posed by sparse data.
In the context of clinical applications, Heller et al. [153] provide the kits19 dataset, which includes 300 kidney tumor cases with CT semantic segmentations and surgical outcomes, aiming to enhance research in KC by making rich datasets available. This effort underscores the critical role of accessible high-quality datasets in advancing early tumor detection and improving surgical outcomes. Flitcroft et al. [154] and Nassour et al. [155] both discuss the scarcity of reliable diagnostic biomarkers for KC, particularly in non-invasive strategies like urinary protein analysis and the evaluation of Lynch syndrome’s impact on KC. These studies highlight the gap in epidemiological data and the need for comprehensive datasets to develop effective diagnostic tools.
Technological advancements in handling data scarcity are also evident in DL applications. Alzu’bi et al. [156] and Bhattacharjee et al. [157] discuss the use of CNNs for kidney tumor classification and detection. They focus on optimizing DL approaches to perform effectively even with limited data, a crucial advancement given the rarity and variability of KC data. Lastly, Wang et al. [158] delve into the utilization of DL techniques for the imaging diagnosis of RCC. They emphasize how these advanced computational tools allow clinicians to diagnose and evaluate renal tumors more accurately and swiftly. One of the key strategies discussed is leveraging unlabeled data, which is a valuable solution for coping with data scarcity. By utilizing semi-supervised learning techniques, the potential of existing datasets can be maximized without the need for extensive labeled data.
6.3 Data bias and generalization
@@
6.4 Privacy concerns, expertise shortage and accuracy instability
@@
6.5 Knowledge gain and data annotation
@@
6.6 Negative transfer and overfitting